1. Show that, for every n > 1: n ka n(n + 1)(2n +1) 6 k=1
Answer #1

n n(n+1)(2041) K=1 FOR L.H.S I= 1(1+1)(2+0 RoHS 6 1. for hold. FOR n=2 L.H.S 5 R. HIS 30 212+1)(4+1) 6 = 5 for n = 2 0 ý hold-->{23 (n+1){ 20+70+6} 11 6840) { 20(*+2)+3(v+2) principal of induction theorem Now (M+1) { 20+ pt 6P+6 6 (+) { 287 49 + 30+6